A man distributes £2 among 20 boys, giving 5p each to some
and 25p to the rest. How many boys received 25p each.

Found 2 solutions by josgarithmetic, TimothyLamb:
Answer by josgarithmetic(39620)

x = how many received 5p
y = how many received 25p

Given x+y=20 boys

Account for the 2 pound value, 5x+25y=2

System to solve is:
-------------------------
5x+25y=2
-------------------------
x+y=20
-------------------------

Solve the system. Still need help?

Answer by TimothyLamb(4379)

---
x + y = 20
5x + 25y = 200
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 15 boys got 5p
y= 5 boys got 25p
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php